Problem 4. Find the derivative of the faction yan - VI Problems. Find the derivative of the function y=sin( 2.1) Problem 6.
Answer #1

Problems to f(x)= x²+x²_x fca) = 3x +22. put f(x) = 0 3*2 + 2x = 0 2 (3x+2) = 0 2-O 2 - Critical points. 2- -2Problem feos = x²6x²+5 ; E3,5] fied absolute maaima and minima Sol fw 3x² _ 128 Put f(x) = 0 372-122=0 X(32-12) 220 x=4 Now,when x= 4, f(x) > 0 absolute minimaat 2 Pooblem 9 local maxima (2,01 continous but not differentiable at x = 2 Problem 8 X C) (2,0) 3,0) (4,0)(50)